# The Impact of Imperfect XAI on Human-AI Decision-Making

### v3

@@ 4.4. Explanation Modalities
Following previous studies (Winter et al., 2015; Calisto et al., 2023; Pacheco and Martinho, 2019), we define assertive explanations to include words and adjectives such as “definitely” and “clearly”. We define the non-assertive explanations to include words and adjectives such as “might be” and “appears to be”. For the neutral condition, we omit the adjectives to maintain the same structure of the information being presented. For the natural language explanations, we append the assertiveness to the beginning of the explanation generated by the model to read like a sentence. For the non-assertive and assertive conditions, we removed the text “this is a” from the generated explanation in order to incorporate it into the sentence structure we designed. The three versions of assertiveness for both explanation modalities are shown in Figure 6 on p. 6. To our knowledge, there is no literature to rationalize how to appropriately present assertiveness visually for example-based explanations, so we opt to use natural language in combination with the example-based explanations as seen in Figure 6.
@@ 4.5. Recruitment
We recruit the participants through several communication channels that are related to the environment and conservation, such as the AI for Conservation Slack, Birding International Discord, Climate Change AI community forum, WildLabs.net community forum, and Audubon Society mailing lists. Additionally, we use Prolific as previous research has indicated that this platform is a reliable source of research data (Peer et al., 2017; Palan and Schitter, 2018). We apply a custom filter on Prolific to target individuals who currently work in a field related to nature, science, the environment, or animals. Overall, we try to limit recruitment to only address people with prior knowledge of birding to minimize the prevalence of novices’ randomly guessing bird species identification. All participants are required to live in the United States, Canada, or Europe, be over 18 years of age, and be fluent in English. Based on informal conversations with birding experts from the Birding International Discord, we decide to include participants from Canada and Europe because bird experts can be familiar with bird species in North America regardless of where they live. Eligible participants that complete the entire study are compensated with a 5 USD payment or an Amazon gift card, equivalent to a pay of 12.5 USD per hour on average.
After excluding participants that provide incomplete and fake responses (i.e., lorem ipsum response to our survey question), we have $136$ people complete our study. In order to determine if a participant is familiar with birding, participants take a bird identification test (phase A in Figure 3). The participant’s score on the bird identification test is used to determine whether they are considered a non-expert or an expert. Details related to clustering participants based on their test scores are provided in Section 5.
@@ 4.6. Quantitative and Qualitative Metrics
We quantitatively calculate appropriate reliance across the four dimensions defined by Schemmer et al. (2023): correct AI reliance, correct self-reliance, under-reliance, and over-reliance. Accordingly, correct AI reliance measures the number of correct decisions when the human’s initial decision is incorrect, and the human is rightly taking over the correct AI advice. Correct self-reliance is when the human initially makes the correct decision and does not overwrite their decision with the incorrect AI advice. On the other hand, under-reliance reflects the case in which the human initially makes an incorrect decision and does not adhere to the correct AI advice. Contrarily, over-reliance represents the scenario in which the human makes an initial correct decision but overrides her own decision with the incorrect AI advice. Following the appropriate reliance metrics defined by  Schemmer et al. (2023), we calculate RSR and RAIR to account for the appropriateness of reliance.
@@
According to Schemmer et al.’s conceptualization of Appropriateness of Reliance (Schemmer et al., 2023), this results in the following:
@@
Lastly, as defined by previous work (e.g.,  (Hemmer et al., 2021; Bansal et al., 2021; Green and Chen, 2019)), we can calculate the human-AI team performance to determine if CTP exists. Following the constructs defined in those previous works, we determine if CTP exists by calculating the participants’ performance in identifying the bird species before and after they see the AI advice and compare this to the performance of the model on the twelve birds images shown to the participant. We utilize accuracy as performance metric. Since every participant is shown six birds that the AI correctly classifies and six that the AI incorrectly classifies, the model performance is $50\%$.
@@ 5. Results
To answer our research questions and confirm or reject our hypotheses, we conduct rigorous statistical and qualitative analyses. We measure appropriate reliance based on metrics defined in previous work: Relative AI reliance (RAIR) and relative self-reliance (RSR) (Schemmer et al., 2023). By doing so, we answer RQ 1 and RQ 2 in Section 5.2. We also calculate the participant’s task accuracy before and after receiving AI and XAI advice to answer RQ 3 (see Section 5.3). This way, we can determine whether complementary team performance exists (Hemmer et al., 2021). Additionally, based on the new metric, Deception of Reliance (DoR), we measure to what extent explanations deceive humans, answering RQ 4 in Section 5.4. Lastly, we qualitatively analyze open-ended responses through rigorous inductive content analysis in Section 5.5. For all of our research questions, we look at two different types of explanations: natural language explanations that are focused on specific features present in the image and visual, example-based explanations showing the top three most similar example images from the training set. While our analyses look at both modalities, we do not intend to directly compare them. Therefore, we do not conclude one modality as better or worse than the other.
@@ 5.2. Moderating Effects in Imperfect XAI Research Model
assertiveness
Z1
Z2
neutral
0
0
non-assertive
1
0
assertive
0
1
@@
coeff
p
coeff
p
coeff
p
coeff
p
const
1.26
.00
.43
.17
-17.16
.98
-3.60
.00
corr
.57
.29
1.02
.04
13.25
.98
-.4.36
.74
exp
2.12
.00
-1.25
.00
15.89
.98
3.25
.00
Z1
-.46
.24
.26
.47
.14
.26
-.09
.83
Z2
.00
1.00
.00
1.00
-.31
.58
.09
.83
exp x corr
-1.00
.05
-1.04
.03
-13.33
.98
-.73
.57
Z1 x corr
.46
.44
-.03
.95
-.28
.71
-.45
.55
Z2 x corr
.19
.74
-.16
.77
.90
.23
-.43
.55
@@ 5.2.3. Participant’s level of expertise has a direct effect on RSR for natural language explanations
In addition to analyzing whether the level of expertise and assertiveness moderate the effect of explanations’ correctness on RAIR, we conduct the same analyses for the effect of explanations’ correctness on RSR. For RSR, we look at all cases in which the AI prediction is giving incorrect advice (i.e., the prediction is wrong) and the initial human decision is correct (Schemmer et al., 2023).
@@
Thus, we perform a regression analysis with the level of expertise and assertiveness as independent variables and drop the interaction terms. We observe that there is no significant effect of assertiveness on RSR (coeff =.$=.10$, p-value =.$=.58$). However, the level of expertise (coeff $=3.18$, p-value =.$=.00$) has a significant effect on RSR. With a positive coefficient, this means that experts dismiss incorrect AI advice more than non-experts when shown natural language explanations. This can also be seen in Figure 8, which tells us that experts have a higher RSR than non-experts. Therefore, we reject hypothesis 2, and additionally, hypothesis 4 as the level of expertise does not take in a moderating role but has a direct effect on RSR for natural language explanations. On top of that, we reject hypothesis 6.
@@
We can see in Table 1 that the level of expertise does not significantly moderate the effect of correctness on RSR (coeff = -.73, p-value = .57). Additionally, there is no significant moderation of assertiveness (Z1 x corr.: coeff =.$=-.45$, p-value =.$=.55$; Z2 x corr.: coeff =.$=-.43$, p-value =.$=.55$). Thus, we conduct a regression analysis without the interaction terms. The results of this analysis show no direct effect of assertiveness on RSR (coeff =.$=-.03$, p-value =.$=.86$). However, there is a significant direct effect of explanations’ correctness on RSR (coeff $=-1.40$, p-value =.$=.00$) and a direct effect of level of expertise on RSR (coeff $=3.05$, p-value $=0.00$). This means that experts more often correctly override the wrong AI advice and stick to their correct initial decision compared to non-experts for example-based explanations. Additionally, when the explanations are correct, participants more often correctly override wrong AI advice and stick to their correct initial decision. Hence, experts have a higher RSR than non-experts which can also be seen in Figure 8. Moreover, as incorrect explanations have a higher impact on RSR, experts are able to better identify false AI advice for incorrect explanations. This also shows that experts are able to identify incorrect AI advice to a greater extent than non-experts; experts, in this case, rely more heavily on their own judgment.
@@ 5.3. Human-AI Team Performance
We can separate this figure into correct and incorrect explanations. When we only consider cases with correct explanations (Figure 11 in Section A.4), the non-experts’ task accuracy is approximately the same as the AI alone: $48.81\%$ for natural language explanations and $49.33\%$ for example-based explanations. Experts reach CTP in both modalities. When only considering incorrect explanations (Figure 12 in Section A.4), we still see complementary team performance for the experts. However, the non-experts’ task accuracy suffers more when shown incorrect explanations. Non-experts’ task accuracy for natural language explanations is $42.86\%$ and $36.67\%$ for example-based explanations.
@@ 5.4. Deception caused by Imperfect XAI
The figure shows that experts have a higher RSR than non-experts for both incorrect and correct explanations across both explanation modalities validating that experts rely more on their own initial decision when AI advice is given. The most striking result that emerges from the data is that for example-based explanations, we observe that experts have a significantly higher RSR for incorrect explanations (RSR $=0.57$) than correct explanations (RSR $=0.29$), resulting in a negative $DoR_{RSR}$ of $-0.28$ (p-value $=0.00$). As a result, experts are falsely relying on the AI advice when provided with correct example-based explanations [^7] Note that correct example-based explanations are consistent in showing three images of the predicted class. Incorrect example-based explanations represent three images that do not correspond to the predicted class of the AI. Moreover, the examples shown are not consistent with the bird species displayed in 90% of the correct advice, incorrect explanation cases and in 40% of the incorrect advice, incorrect explanation cases in our study.. This means that experts are prone to being misled by correct explanations when the AI advice is incorrect. However, we do not see this trend for natural language explanations. Here, there is a positive $DoR_{RSR}$ of 0.09, which is not significant (p-value $=0.41$). For both modalities, the $DoR_{RAIR}$ is positive, meaning that experts rightly follow correct AI advice more often when provided with correct explanations than with incorrect explanations. The data shows a weak significant positive deception of reliance for example-based explanations ($DoR_{RAIR}=0.16$, p-value $=0.10$). Similarly to the RSR cases, for the RAIR cases, the experts are being provided with three consistent examples for correct explanations that represent the AI’s correctly predicted bird species. The incorrectly provided explanations represent three images that can be inconsistent in the bird species. Thus, experts are deceived by such incorrect explanations even though the AI advice is correct.
@@ 5.5. Designing for Imperfect XAI
At the end of the bird identification task, we ask participants: “Under what circumstances would you prefer assertive (e.g., “definitely”, “clearly”) versus non-assertive (e.g., “might be”, “appears to be”) versus neutral explanations and why?”. With imperfect XAI existing in human-AI collaborations, it is necessary to not only understand quantitatively how it impacts decision-makers but also qualitatively. Even though we do not observe the level of assertiveness to have a direct or moderation effect on appropriate reliance, it is still valuable to analyze participants’ preferences when it comes to the tone of explanations.
@@ 5.5.1. Aggregated Dimension: AI Behaviors
Considering individual predictions on a more granular level, most participants ($31$ out of $136$) make comments related to the confidence of the AI’s prediction. These participants comment on how this factor could be used to determine the tone of the explanation. Specifically, participants, “… would prefer the level of assertiveness to depend on the level of confidence of the answer given by the AI”. One participant expands upon that sentiment by specifying when non-assertive versus assertive tones should be used: “I would prefer assertive sentences when the probability of the AI model is very high, while I would prefer non-assertive when the probability is very close to other classes of the model”.
$9$ out of $136$ participants make comments related to the correctness and quality of the AI explanations. One participant who sees example-based explanations comments on how some of the examples are incorrect and do not show the correct species. They use this specific situation to rationalize when they would prefer the tone of explanations to be assertive versus non-assertive: “I would prefer assertive explanations if the ”similar” photos were actually of the correct species and if the explanation confirmed this. Otherwise, non-assertive explanations are more helpful”. Another participant that makes a comment related to this theme agrees that assertive language should be used, “when all of the reference pictures match up and there are no other similar-looking species”.
@@ 5.5.2. Aggregated Dimension: Impact on Human-AI Team
$15$ out of the $136$ make comments related to the impacts on the decision-maker. Some comments consist of concerns related to being misled and over- or under-relying on the AI, while other comments motivate the benefits of having assertive explanations. For example, one participant states that “Assertive words create more security while changing your opinion or trying to gain knowledge”. Another participant that shares the same sentiment said “I would prefer assertive explanations because it would make me feel more secure about the answer”. However, some participants do not share the same sentiment about assertive explanations and pointed out the potential consequences of them: “Assertive AI explanations were given for incorrect identifications, which would mislead users”. Given that potential to be misled, another participant rationalizes they “… would prefer non-assertive explanations because I [they] do not fully trust AI with bird ID just yet”.
$17$ out of the $136$ participants rationalize that the assertiveness of explanations should be based on how ambiguous the input is for a given prediction. For example, one participant brings up the quality of the input image and the difficulty of the bird ID as a way to determine whether explanations should be assertive or not: “When it comes to less distinctive ID’s like most sparrows, or harder to ID circumstances like winter or females or juveniles, or situations with weird lighting or harder angles it makes sense to use non-assertive.”. On a similar line of thought, another participant says, “I would prefer assertive explanations for birds that have distinctive traits over similar bird species, and non-assertive or neutral explanations for birds that are similar with characteristics that are more difficult to tell apart

### v4

@@ 4.4. Explanation Modalities
Following previous studies (Winter et al., 2015; Calisto et al., 2023; Pacheco and Martinho, 2019), we define assertive explanations to include words and adjectives such as “definitely” and “clearly”. We define the non-assertive explanations to include words and adjectives such as “might be” and “appears to be”. For the neutral condition, we omit the adjectives to maintain the same structure of the information being presented. For the natural language explanations, we append the assertiveness to the beginning of the explanation generated by the model to read like a sentence. For the non-assertive and assertive conditions, we removed the text “this is a” from the generated explanation in order to incorporate it into the sentence structure we designed. The three versions of assertiveness for both explanation modalities are shown in Figure 6. To our knowledge, there is no literature to rationalize how to appropriately present assertiveness visually for example-based explanations, so we opt to use natural language in combination with the example-based explanations.
@@ 4.5. Recruitment
We recruit the participants through several communication channels that are related to the environment and conservation, such as the AI for Conservation Slack, Birding International Discord, Climate Change AI community forum, WildLabs.net community forum, and Audubon Society mailing lists. Additionally, we use Prolific as previous research has indicated that this platform is a reliable source of research data (Peer et al., 2017; Palan and Schitter, 2018). We apply a custom filter on Prolific to target individuals who currently work in a field related to nature, science, the environment, or animals. Participants receive compensation that is above minimum wage. Overall, we try to limit recruitment to only address people with prior knowledge of birding to minimize the prevalence of novices’ randomly guessing bird species identification. After excluding participants who provide incomplete and fake responses (i.e., lorem ipsum response to our survey question), we have $136$ people complete our study. In order to determine if a participant is familiar with birding, participants take a bird identification test (phase A in Figure 3). The participant’s score on the bird identification test is used to determine whether they are considered a non-expert or an expert. Details related to clustering participants based on their test scores are provided in Section 5.
@@ 4.6. Quantitative and Qualitative Metrics
We quantitatively calculate appropriate reliance across the four dimensions defined by Schemmer et al. (2023b): correct AI reliance, correct self-reliance, under-reliance, and over-reliance. Accordingly, correct AI reliance measures the number of correct decisions when the human’s initial decision is incorrect, and the human is rightly taking over the correct AI advice. Correct self-reliance is when the human initially makes the correct decision and does not overwrite their decision with the incorrect AI advice. On the other hand, under-reliance reflects the case in which the human initially makes an incorrect decision and does not adhere to the correct AI advice. On the other hand, over-reliance represents the scenario in which the human makes an initial correct decision but overrides her own decision with incorrect AI advice. Following the appropriate reliance metrics defined by  Schemmer et al. (2023b), we calculate RSR and RAIR to account for the appropriateness of reliance.
@@
According to Schemmer et al.’s conceptualization of Appropriateness of Reliance (Schemmer et al., 2023b), this results in the following:
@@
Lastly, as defined by previous work (e.g.,  (Hemmer et al., 2021; Bansal et al., 2021; Green and Chen, 2019)), we can calculate the human-AI team performance to determine if CTP exists. Following the constructs defined in those previous works, we determine if CTP exists by calculating the participants’ performance in identifying the bird species before and after they see the AI advice and compare this to the performance of the model on the twelve birds images shown to the participant. We utilize accuracy as a performance metric. Since every participant is shown six birds that the AI correctly classifies and six that the AI incorrectly classifies, the model performance is $50\%$.
@@ 5. Results
To answer our research questions and confirm or reject our hypotheses, we conduct rigorous statistical and qualitative analyses. We measure appropriate reliance based on metrics defined in previous work: Relative AI reliance (RAIR) and relative self-reliance (RSR) (Schemmer et al., 2023b). By doing so, we answer RQ 1 and RQ 2 in Section 5.2. We also calculate the participant’s task accuracy before and after receiving AI and XAI advice to answer RQ 3 (see Section 5.3). This way, we can determine whether complementary team performance exists (Hemmer et al., 2021). Additionally, based on the new metric, Deception of Reliance (DoR), we measure to what extent explanations deceive humans, answering RQ 4 in Section 5.4. Lastly, we qualitatively analyze open-ended responses through rigorous inductive content analysis in Section 5.5. For all of our research questions, we look at two different types of explanations: natural language explanations that are focused on specific features present in the image and visual, example-based explanations showing the top three most similar example images from the training set. While our analyses look at both modalities, we do not intend to compare them directly. Therefore, we do not conclude one modality is better or worse than the other.
@@ 5.2. Moderating Effects in Imperfect XAI Research Model
assertiveness Z1 Z2
neutral 0 0
non-assertive 1 0
assertive 0 1
@@
coeff p coeff p coeff p coeff p
const 1.26 .00 .43 .17 -17.16 .98 -3.60 .00
corr .57 .29 1.02 .04 13.25 .98 -.4.36 .74
exp 2.12 .00 -1.25 .00 15.89 .98 3.25 .00
Z1 -.46 .24 .26 .47 .14 .26 -.09 .83
Z2 .00 1.00 .00 1.00 -.31 .58 .09 .83
exp x corr -1.00 .05 -1.04 .03 -13.33 .98 -.73 .57
Z1 x corr .46 .44 -.03 .95 -.28 .71 -.45 .55
Z2 x corr .19 .74 -.16 .77 .90 .23 -.43 .55
  • 1
    corrcorrectness; explevel of expertise
@@ 5.2.3. Participant’s level of expertise has a direct effect on RSR for natural language explanations
In addition to analyzing whether the level of expertise and assertiveness moderate the effect of explanations’ correctness on RAIR, we conduct the same analyses for the effect of explanations’ correctness on RSR. For RSR, we look at all cases in which the AI prediction is giving incorrect advice (i.e., the prediction is wrong) and the initial human decision is correct (Schemmer et al., 2023b).
@@
Thus, we perform a regression analysis with the level of expertise and assertiveness as independent variables and drop the interaction terms. We observe that there is no significant effect of assertiveness on RSR (coeff =.$=.10$, p-value =.$=.58$). However, the level of expertise (coeff $=3.18$, p-value =.$=.00$) has a significant effect on RSR. With a positive coefficient, this means that experts dismiss incorrect AI advice more than non-experts when shown natural language explanations. This can also be seen in Figure 8, which tells us that experts have a higher RSR than non-experts. Therefore, we reject hypothesis 2, and additionally, hypothesis 4 as the level of expertise does not have a moderating role but has a direct effect on RSR for natural language explanations. On top of that, we reject hypothesis 6.
@@
We can see in Table 1 that the level of expertise does not significantly moderate the effect of correctness on RSR (coeff = -.73, p-value = .57). Additionally, there is no significant moderation of assertiveness (Z1 x corr.: coeff =.$=-.45$, p-value =.$=.55$; Z2 x corr.: coeff =.$=-.43$, p-value =.$=.55$). Thus, we conduct a regression analysis without the interaction terms. The results of this analysis show no direct effect of assertiveness on RSR (coeff =.$=-.03$, p-value =.$=.86$). However, there is a significant direct effect of explanations’ correctness on RSR (coeff $=-1.40$, p-value =.$=.00$) and a direct effect of level of expertise on RSR (coeff $=3.05$, p-value $=0.00$). This means that experts more often correctly override the wrong AI advice and stick to their correct initial decision compared to non-experts for example-based explanations. Additionally, when the explanations are correct, participants more often correctly override wrong AI advice and stick to their correct initial decision. Hence, experts have a higher RSR than non-experts, which can also be seen in Figure 8. Moreover, as incorrect explanations have a higher impact on RSR, experts are able to identify false AI advice for incorrect explanations better. This also shows that experts are able to identify incorrect AI advice to a greater extent than non-experts; experts, in this case, rely more heavily on their own judgment.
@@ 5.3. Human-AI Team Performance
We can separate Figure 7 into correct and incorrect explanations. When we only consider cases with correct explanations (Figure 11 in Section A.4), the non-experts’ task accuracy is approximately the same as the AI alone: $48.81\%$ for natural language explanations and $49.33\%$ for example-based explanations. Experts reach CTP in both modalities. When only considering incorrect explanations (Figure 12 in Section A.4), we still see complementary team performance for the experts. However, the non-experts’ task accuracy suffers more when shown incorrect explanations. Non-experts’ task accuracy for natural language explanations is $42.86\%$ and $36.67\%$ for example-based explanations.
@@ 5.4. Deception caused by Imperfect XAI
The figure shows that experts have a higher RSR than non-experts for both incorrect and correct explanations across both explanation modalities, validating that experts rely more on their own initial decisions when AI advice is given. The most striking result that emerges from the data is that for example-based explanations, we observe that experts have a significantly higher RSR for incorrect explanations (RSR $=0.57$) than correct explanations (RSR $=0.29$), resulting in a negative $DoR_{RSR}$ of $-0.28$ (p-value $=0.00$). As a result, experts are falsely relying on the AI advice when provided with correct example-based explanations [^7] Note that correct example-based explanations are consistent in showing three images of the predicted class. Incorrect example-based explanations represent three images that do not correspond to the predicted class of the AI. Moreover, the examples shown are not consistent with the bird species displayed in 90% of the correct advice, incorrect explanation cases and in 40% of the incorrect advice, incorrect explanation cases in our study.. This means that experts are prone to being misled by correct explanations when the AI advice is incorrect. However, we do not see this trend for natural language explanations. Here, there is a positive $DoR_{RSR}$ of 0.09, which is not significant (p-value $=0.41$). For example-based explanations, the $DoR_{RAIR}$ is positive, meaning that experts rightly follow correct AI advice more often when provided with correct explanations than with incorrect explanations. The data shows a weak, significant positive deception of reliance for example-based explanations ($DoR_{RAIR}=0.16$, p-value $=0.10$). Similarly to the RSR cases, for the RAIR cases, the experts are provided with three consistent examples for correct explanations that represent the AI’s correctly predicted bird species. The incorrectly provided explanations represent three images that can be inconsistent in the bird species. Thus, experts are deceived by such incorrect explanations even though the AI advice is correct.
@@ 5.5. Designing for Imperfect XAI
At the end of the bird identification task, we ask participants: “Under what circumstances would you prefer assertive (e.g., “definitely”, “clearly”) versus non-assertive (e.g., “might be”, “appears to be”) versus neutral explanations and why?”. With imperfect XAI existing in human-AI collaborations, it is necessary not only to understand quantitatively how it impacts decision-makers but also qualitatively. Even though we do not observe the level of assertiveness to have a direct effect or a moderation effect on appropriate reliance, it is still valuable to analyze participants’ preferences when it comes to the tone of explanations.
@@ 5.5.1. Aggregated Dimension: AI Behaviors
Considering individual predictions on a more granular level, many participants ($31$ out of $136$) make comments related to the confidence of the AI’s prediction. These participants comment on how this factor could be used to determine the tone of the explanation. Specifically, participants, “… would prefer the level of assertiveness to depend on the level of confidence of the answer given by the AI”. One participant expands upon that sentiment by specifying when non-assertive versus assertive tones should be used: “I would prefer assertive sentences when the probability of the AI model is very high, while I would prefer non-assertive when the probability is very close to other classes of the model”.
$9$ out of $136$ participants make comments related to the correctness and quality of the AI explanations. One participant who sees example-based explanations comments on how some of the examples are incorrect and do not show the correct species. They use this specific situation to rationalize when they would prefer the tone of explanations to be assertive versus non-assertive: “I would prefer assertive explanations if the ‘similar’ photos were actually of the correct species and if the explanation confirmed this. Otherwise, non-assertive explanations are more helpful”. Another participant who makes a comment related to this theme agrees that assertive language should be used, “when all of the reference pictures match up and there are no other similar-looking species”.
@@ 5.5.2. Aggregated Dimension: Impact on Human-AI Team
$15$ out of the $136$ make comments related to the impacts on the decision-maker. Some comments consist of concerns related to being misled and over- or under-relying on the AI, while other comments motivate the benefits of having assertive explanations. For example, one participant states that “Assertive words create more security while changing your opinion or trying to gain knowledge”. Another participant who shares the same sentiment said, “I would prefer assertive explanations because it would make me feel more secure about the answer”. However, some participants do not share the same sentiment about assertive explanations and pointed out the potential consequences of them: “Assertive AI explanations were given for incorrect identifications, which would mislead users”. Given that potential to be misled, another participant rationalizes they “… would prefer non-assertive explanations because I [they] do not fully trust AI with bird ID just yet”.
$17$ out of the $136$ participants rationalize that the assertiveness of explanations should be based on how ambiguous the input is for a given prediction. For example, one participant brings up the quality of the input image and the difficulty of the bird ID as a way to determine whether explanations should be assertive or not: “When it comes to less distinctive IDs like most sparrows, or harder to ID circumstances like winter or females or juveniles, or situations with weird lighting or harder angles it makes sense to use non-assertive.”. On a similar line of thought, another participant says, “I would prefer assertive explanations for birds that have distinctive traits over similar bird species, and non-assertive or neutral explanations for birds that are similar with characteristics that are more difficult to tell apart”.